OpenStudy (kwonbon):

If sin(x) = 1/3 and sec(y) = 25/24, where x and y lie between 0 and π/2, evaluate the expression using trigonometric identities. sin(x + y)

jhonyy9 (jhonyy9):

do you know formula sin(a+b) = ?

jhonyy9 (jhonyy9):

@OtherWorldly

OpenStudy (kwonbon):

Yes sin A cos B + cos A sin B ? @jhonyy9

jhonyy9 (jhonyy9):

and sec(y) = ?

OpenStudy (kwonbon):

1/cos

OpenStudy (kwonbon):

@jackthegreatest

jhonyy9 (jhonyy9):

ok evaluate using this values the sin(x+y) formula

jhonyy9 (jhonyy9):

1/cos = 25/24 -- given than cos = ?

OpenStudy (kwonbon):

ummm. 24/25? so would the formula then be: sin(1/3)cos(24/25) + sin(24/25)cos(1/3) ?

jhonyy9 (jhonyy9):

you not are sure in this ? why ?

OpenStudy (kwonbon):

um... what? LOL

OpenStudy (kwonbon):

this would be in radians, correct?

OpenStudy (jackthegreatest):

@jhonyy9 the sum and difference formula?

OpenStudy (jackthegreatest):



OpenStudy (kwonbon):

@jackthegreatest so then, using that formula, would my answer be correct? sin(1/3)cos(24/25) + sin(24/25)cos(1/3)? (in radians)

jhonyy9 (jhonyy9):

how you understand this << sum and difference >> ?

OpenStudy (jackthegreatest):

@jhonyy9 well thats what precalc textbooks call it

jhonyy9 (jhonyy9):

@zepdrix please do you understand this ?

jhonyy9 (jhonyy9):

sin(x+y) so we need using here the sum just and not the difference - why said it so ?

OpenStudy (kwonbon):

i'm so lost omg

OpenStudy (jackthegreatest):

@jhonyy9 btw, those are angles

jhonyy9 (jhonyy9):

@jackthegreatest you wrote it hence above - please clarify

OpenStudy (jackthegreatest):

degrees

jhonyy9 (jhonyy9):

@zepdrix sorry ,do you understand this ,,difference" what mean here ?

OpenStudy (jackthegreatest):

@jhonyy9 what would be best for kwonbon is if you just worked out the problem

jhonyy9 (jhonyy9):

we need just giving the way ,help the askers understanding how he or she can solving the posted exercise and not solve from start till the end this exercise,problem or question hope you know it sure

jhonyy9 (jhonyy9):

@zepdrix do you agree this please ?

OpenStudy (kwonbon):

@jhonyy9 but you're not even helping me LOL it feels like you don't even know how to do the problem yourself...

OpenStudy (kwonbon):

like you don't even answer the questions i ask in response to your questions that you ask me

jhonyy9 (jhonyy9):

than you check there above i drived you step by step - yes ? and o ve got the formula sin(x+y) = ?

OpenStudy (kwonbon):

._. yes

OpenStudy (kwonbon):

so then i asked you if it was sin(1/3)cos(24/25) + sin(24/25)cos(1/3)

jhonyy9 (jhonyy9):

no bc. given that sin x = 1/3 or cos x = 24/25

zepdrix (zepdrix):

Oh sorry I ran off :3

jhonyy9 (jhonyy9):

there you need substituting the sinx by 1/3 not sin1/3 and same the cos x

zepdrix (zepdrix):

So we agreed that the Angle Addition Formula would be useful here,\[\large\rm \sin(x+y)=\color{orangered}{\sin x} \cos y+\sin y \cos x\]Look at this thing in orange a sec. They gave us this information at the very start.\[\large\rm \color{orangered}{\sin x=\frac13}\]This is telling us that we can replace sinx with 1/3 in our expansion,\[\large\rm \sin(x+y)=\color{orangered}{\frac13} \cos y+\sin y \cos x\]Kwon, does that help you to understand why your answer was nonsense? There shouldn't be a sin in front of your 1/3 anymore.

zepdrix (zepdrix):

Oh he left :( ugh

jhonyy9 (jhonyy9):

thank you zepdrix

jhonyy9 (jhonyy9):

@jackthegreatest do you understand it now ?

OpenStudy (jackthegreatest):

what you mean if i understand it now?

jhonyy9 (jhonyy9):

how you get the right answer ?

jhonyy9 (jhonyy9):

when evaluate the sin(x+y) so there you need writing not sinx just 1/3 bc. sin x =1/3 and cos x = 24/25

jhonyy9 (jhonyy9):

@kwonbon

OpenStudy (jackthegreatest):

haha you realize all i was saying is that x and y are degrees

jhonyy9 (jhonyy9):

what i dont understand that why wrote there above the sum and difference when here need using just the sum ?

OpenStudy (sshayer):

\[\sin \left( x+y \right)=\sin x \cos y+\cos x \sin y\] \[\sin x=\frac{ 1 }{ 3 },\cos x=\sqrt{1-\sin ^2x}=\sqrt{1-\frac{ 1 }{ 9 }}=\sqrt{\frac{ 8 }{ 9 }}=\frac{ 2\sqrt{2} }{ 3 }\] \[\cos y=\frac{ 24 }{ 25 },\sin y=\sqrt{1-\frac{ 576 }{ 625 }}=\sqrt{\frac{ 49 }{ 625 }}=\frac{ 7 }{ 25 }\]

OpenStudy (sshayer):

plug the values in the formula and find sin(x+y)

OpenStudy (mathmale):

Have a look at "sin(1/3)cos(24/25) + sin(24/25)cos(1/3)." You are told that sin x =1/3 and sec y=24/25. The sum formula for the sine function requires sin x, cos y, sin y and cos x. Where are you going to get these values for these four? The value of sin x is given: sin x = 1/3. Here, x is an angle and 1/3 is the value of the sine function. So writing sin(1/3) makes no sense. What you want to do is to replace sin x in the sum formula with 1/3. The value of sec y is given: 25/24. Taking the reciprocal of both sides, cos y=24/25. Do not write cos (24/25); instead, substitute 24/25 for cos y. Can you find sin y? Can you find cos x? Can you now complete the problem?

OpenStudy (mathmale):

Acknowledging that sshayer has already found the values of all four necessary trig functions.